OpenStudy (anonymous):

Solve the rational inequality. Write the solution set in interval notation. x+9 ---- < 7 x+4

OpenStudy (anonymous):

pain in the butt you have to start with \[\frac{x+9}{x+4}-7<0\]

OpenStudy (anonymous):

then get \[\frac{x+9-7(x+4)}{x+4}<0\] \[\frac{-6x-19}{x+4}<0\]

OpenStudy (anonymous):

zeros are at -4 and \[\frac{19}{6}\] and it will be positive between them and negative outside of them so your answer is \[x<-4\] or \[x>\frac{19}{6}\]